OpenStudy (tdskitz):

How many solutions does the following equation have? |4x + 12| = 0 A) No Solution B) One Solution C) Two Solutions D) Infinetly Many Solutions

OpenStudy (kamek):

oh well that is easy: This has one solution

OpenStudy (icrazyaliens):

It has two solutions... need me to explain?

OpenStudy (mathmale):

Hint: y=|x| as a v-shaped graph, with its vertex at (0,0). Your |4x + 12| = 0 could be factored: 4|x+3| = 0. Since 4 never equals 0, focus on solving |x+3|=0. How many solutions would this have? Suggestion: graph y=|x+3|. Start with the graph of y=|x|, and then translate this graph horizontally as appropriate.